# A subcopula characterization of dependence for the Multivariate Bernoulli Distribution

### v1

@@ 4 Trivariate Bernoulli
where (42) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $10<\theta<$ 1 then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence. In fact, in terms of the probabilities $p_{ijk}$ this implies $p_{000}=\theta,$ $p_{111}=1-\theta$ and zero probability for the remaining six cases.
@@ 5 Multivariate Bernoulli
Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ Accordingly to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
$S^{(n)}(u_{1},\ldots,u_{n})=\begin{cases}\,1&\mbox{ if }u_{r}=1\mbox{ for all }r\in\{1,\ldots,n\},\\ \,\theta_{s}&\mbox{ if }u_{s}=\theta_{s}\mbox{ and }u_{r}=1\mbox{ for all }r\neq s,\\ \,\theta_{rs}&\mbox{ if }r<s,u_{r}=\theta_{r},u_{s}=\theta_{s},\mbox{ and }u_{t}=1\mbox{ for all }t\neq r,s,\\ \,\theta_{rst}&\mbox{ if }r<s<t,u_{r}=\theta_{r},u_{s}=\theta_{s},u_{t}=\theta_{t},u_{\ell}=1,\ell\neq r,s,t,\\ \quad\vdots&\qquad\vdots\\ \,\theta_{12\cdots n}&\mbox{ if }u_{1}=\theta_{1},\ldots,u_{n}=\theta_{n},\\ \,0&\mbox{ elsewhere.}\end{cases}$ (43)
with a total of $12^{n}-$ 1 parameters. For any ordered subset of $\{1,\ldots,n\},$ say $A=\{a_{1},\ldots,a_{m}\}$ where $m\leq n$ and if $r<s$ then $a_{r}<a_{s},$ the $m$-variate dependence parameter $\theta_{\!A}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
@@
where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $12^{n}-$ 1 of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $12^{n}-$ 1 . Then expressions for the $\theta_{\!A}$ parameters in terms of (45) probabilities are obtained by:
@@
Algorithm 2
  • Step 0:
    Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
  • Step 1:
    Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the $(n-1)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step 2:
    Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the $(n-2)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step $k$:
[…]
  • Step $n$:
[…]
@@
where from (44) $-1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
@@ 5.1 Bayesian inference
If $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$ is a size $m$ observed random sample from a Multivariate Bernoulli random vector $\mathbf{X}=(X_{1},\ldots,X_{n})$ where $\mathbf{x}_{j}=(x_{j,1},\ldots,x_{j,n})$ for $j\in\{1,\ldots,m\},$ we can make statistical inferences about the $2^{n}$ parameters $p_{i_{1}\ldots i_{n}}.$ For a more simplified notation, following [17], since $(i_{1},\ldots,i_{n})\in\{0,1\}^{n}$ we may consider it as a $n$-positions binary number and apply the one-to-one mapping conversion to a positive integer. For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of $r-1$, so that:
$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$ (48)
@@
[…]
Then the likelihood function associated to the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
[…]
@@
[…]
Clearly (50) has as conjugate prior distribution a $12^{n}-$ 1 dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [1], and therefore the posterior distribution is given by:
@@
5.2 Computational implementation
The Julia programming language [2] has been used for the implementation of all the previous results, and the code is available in a public repository at [4] for full reproducibility. The main functions implemented:
@@ 5.2 Computational implementation
Example 1 revisited. If we choose $\theta=0.6$ then according to (37) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence. Let’s use $\theta_{123}=0.15$ and applying formulas (19) through (26) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ With such values me define in Julia the vector:
julia> pp = [0.15,0.21,0.21,0.03,0.21,0.03,0.03,0.13];
julia> X = MBerDep(pp);
And then verify the we get the expected dependence parameters and measures:
julia> [X.dparam.idx X.dparam.value]
72 Matrix{Any}:
 [1, 2, 3]  0.15
 [1, 2]     0.36
 [1, 3]     0.36
 [2, 3]     0.36
 [1]        0.6
 [2]        0.6
 [3]        0.6
from where we get that the bivariate dependence parameters $\theta_{12}=0.36=\theta_{13}=\theta_{23}=\theta^{2}$ as expected (pairwise independence) but certainly not joint independence since $\theta_{123}=0.15\neq 0.216=\theta^{3}.$ For the dependence measures we get:
julia> [X.dmeas.idx X.dmeas.value]
42 Matrix{Any}:
 [1, 2, 3]  -0.305556
 [1, 2]      0.0
 [1, 3]      0.0
 [2, 3]      0.0
@@ 42 Matrix{Any}:
with pairwise dependencies $\mu_{ij}$ all equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).

### v2

@@ 4 Trivariate Bernoulli
where (32) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $0<\theta<1$ then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence. In fact, in terms of the probabilities $p_{ijk}$ this implies $p_{000}=\theta,$ $p_{111}=1-\theta$ and zero probability for the remaining six cases.
@@ 5 Multivariate Bernoulli
Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ According to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
@@
with a total of $2^{n}-1$ parameters. For any ordered subset of $\{1,\ldots,n\},$ say $A=\{a_{1},\ldots,a_{m}\}$ where $m\leq n$ and if $r<s$ then $a_{r}<a_{s},$ the $m$-variate dependence parameter $\theta_{\!A}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
$\max\left\{\sum_{a\,\in\,A}\theta_{a}-m+1\,,\,0\right\}\,\leq\,\theta_{A}\,\leq\,\min\{\theta_{a}:a\in A\}\,.$ (33)
@@
where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $2^{n}-1$ of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $2^{n}-1.$ Then expressions for the $\theta_{\!A}$ parameters in terms of (34) probabilities are obtained by:
@@
Parameter Selection Procedure 2
  • Step 1: Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
  • Step 2: Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the (n1)(n-1)-variate parameters $\theta_{\!A}$ according to (35).
  • Step 3: Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the (n2)(n-2)-variate parameters $\theta_{\!A}$ according to (35).
  • Step $k+1$: Choose $B_{k}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=k\}$ such that $\sum\bigcup_{t\,=\,0}^{k}B_{t}\leq 1,$ and set the (nk)(n-k)-variate parameters $\theta_{\!A}$ according to (35).
  • Step $n+1$: Calculate $p_{1\cdots 1}=1-\sum\bigcup_{t\,=\,0}^{n-1}B_{t}\,.$
@@
where from (33) $1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
@@ 5.1 Bayesian inference
Under the Bayesian paradigm of statistics, given an observed random sample $\mathbf{y}=(y_{1},\ldots,y_{m})$ from a probability density model $f_{Y}(y\,|\,\theta),$ where $\theta\in\Theta$ is a vector of unknown parameters, any statistical inference about $\theta$ is performed through a posterior distribution for $\theta$ given the data $\mathbf{y}$ applying Bayes’s rule, that is:
$\pi(\theta\,|\,\mathbf{y})=\frac{f(\mathbf{y}\,|\,\theta)\pi(\theta)}{\int_{\,\,\,\Theta}\!\!\!\!\!\!\cdots\!\int f(\mathbf{y}\,|\,\tilde{\theta})\pi(\tilde{\theta})\,d\tilde{\theta}}$
where $\pi(\theta)$ is a prior distribution, and where:
[…]
is also known as the likelihood function. Whenever possible, it is quite convenient to choose as a prior distribution a conjugate family, such that the posterior belongs to that same family, and therefore it is straightforward to identify, since:
$\pi(\theta\,|\,\mathbf{y})\,\propto\,L(\theta\,|\,\mathbf{y})\pi(\theta)\,.$
If $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$ is a size $m$ observed random sample from a Multivariate Bernoulli random vector $\mathbf{X}=(X_{1},\ldots,X_{n})$ where $\mathbf{x}_{j}=(x_{j,1},\ldots,x_{j,n})$ for $j\in\{1,\ldots,m\},$ we can make statistical inferences about the $2^{n}$ parameters $p_{i_{1}\ldots i_{n}}.$ For a more simplified notation, following [1], since $(i_{1},\ldots,i_{n})\in\{0,1\}^{n}$ we may consider it as a $n$-positions binary number and apply the one-to-one mapping conversion to a positive integer. For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of r1r-1, so that:
$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$
@@
$p_{2^{n}}=\mathbb{P}[X_{1}=1,X_{2}=1,\ldots,X_{n}=1]=1-\sum_{r\,=\,1}^{2^{n}-1}p_{r}\,.$
Then the likelihood function associated with the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
$\mathbb{L}(p_{1},\ldots,p_{2^{n}-1}\,|\,\mathbf{x}_{1},\ldots,\mathbf{x}_{m})=\left(\prod_{r\,=\,1}^{2^{n}-1}p_{r}^{\,n_{r}}\right)\left(1-\sum_{r\,=\,1}^{2^{n}-1}p_{r}\right)^{m-\sum_{r\,=\,1}^{2^{n}-1}n_{r}}$ (37)
@@
$n_{r}=\sum_{j\,=\,1}^{m}\mathbf{1}_{\{\mathbf{x}_{j}\,=\,B(r-1)\}}\,.$
Clearly (37) has as conjugate prior distribution a $2^{n}-1$ dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [17], and therefore the posterior distribution is given by:
@@
6 Computational implementation
The Julia programming language [7] has been used for the implementation of all the previous results, and the code is available in a public repository [18] to ensure full reproducibility. Julia offers significant advantages for high-performance numerical computing, especially for high-dimensional matrix operations and intensive statistical computations. Its speed and ability to handle large datasets efficiently make it particularly suitable for the implementation of the proposed methods. Additionally, Julia’s syntax is close to mathematical notation, facilitating the translation of theoretical concepts into computational code, see for example [19]. The main functions implemented:
@@ 6.1 A theoretical example
In Example 1 from Section 4, if we choose $\theta=0.6$ then according to (28) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence. Let’s use $\theta_{123}=0.15$ and applying formulas (12) through (19) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ The Julia code to generate the results from Table 2 is in Appendix A.1, which basically makes use of the MBerDep function described at the beginning of Section 6.
variables dependence parameter dependence measure
$(X_{1},X_{2},X_{3})$ 0.15 -0.3056
$(X_{1},X_{2})$ 0.36 0.0
$(X_{1},X_{3})$ 0.36 0.0
$(X_{2},X_{3})$ 0.36 0.0
Table 2: Dependence parameters and measures corresponding to the provided values in subsection 6.1. The Julia code to generate the results is in Appendix A.1.
@@
All pairwise dependencies $\mu_{ij}$ in Table 2 are equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).
6.2 A simulation example